A swimming pool measuring20.0 m × 12.5 m is filled with water to a depth of 3.75 m. If the initial temperature is 18.4°C, how much heat must be added to the water to raise its temperature to 29.0°C? Assume that the density of water is 1.000 g/mL.
1
Expert's answer
2021-01-25T04:30:06-0500

Given:- "l*w = 20*12.5"

Height = 3.75m

"Volume of pool = l*w*h"

volume=20*12*3.75=937.5"m^3"


To calculate mass we have to use this formula


mass=volume*density
density=1g\ml or 10^6 kg/m^3                                      
mass=937.5*10^6gm

we know that
q=mc(T_2-T_1)
q=937.5*10^6*4.18*(29-18.4) [c=4.18j/g^oc]
q=4.16*10^10j